tyyb:

The length of a rectangular cake pan is 3 more than twice its width. The area of the cake pan is 90 in2. Find the length of the cake pan.

Angle:

Hi @tyyb ! Welcome to QuestionCove! Give me a minute to read your question :)

Angle:

Ok @tyyb , I'm curious because questions like these can be asked in an algebra class is this an algebra question / are you taking algebra?

tyyb:

its a algebra question

Angle:

fantastic so we want to make two algebra equations out of the information given to us "The length of a rectangular cake pan is 3 more than twice its width" Let's make this section: (The length of a rectangular cake pan) = "L" (is) = "=" (3 more than) = "3 + " (twice its width) = "2W" and if we put it all together, our first equation is L = 3 + 2W

Angle:

How about "The area of the cake pan is 90" what equation can be make from that sentence?

Angle:

Our equations are going to be L = 3 + 2W and L * W = 90 do you have any ideas on how we can solve for L or W?